# Robust automatic brain vessel segmentation in 3D CTA scans using dynamic 4D-CTA data

### v1

@@ Abstract
In this study, we develop a novel methodology for annotating the brain vasculature using dynamic 4D-CTA head scans. By using multiple time points from dynamic CTA acquisitions, we subtract bone and soft tissue to enhance the visualization of arteries and veins, reducing the effort required to obtain manual annotations of brain vessels. We then train deep learning models on our ground truth annotations by using the same segmentation for multiple phases from the dynamic 4D-CTA collection, effectively enlarging our dataset by 4 to 5 times and inducing robustness to contrast phases. In total, our dataset comprises 110 training images from 25 patients and 165 test images from 14 patients. In comparison with two similarly-sized datasets for CTA-based brain vessel segmentation, a nnUNet model trained on our dataset can achieve significantly better segmentations across all vascular regions, with an average mDC of 0.846 for arteries and 0.957 for veins in the TopBrain dataset. Furthermore, metrics such as average directed Hausdorff distance (adHD) and topology sensitivity (tSens) reflected similar trends: using our dataset resulted in low error margins (aDHD of 0.304 mm for arteries and 0.078 for veins) and high sensitivity (tSens of 0.877 for arteries and 0.974 for veins), indicating excellent accuracy in capturing vessel morphology. Our code and model weights are available online.
@@ 1 Introduction
Intracranial computed tomographic angiography (CTA) is an essential diagnostic tool that employs intravenous contrast agents to enhance vascular visualization, aiding in the detection of intracranial vascular abnormalities [24]. Since its introduction, CTA has advanced beyond single-phase acquisitions to include multiphase CTA [7], which captures arterial, mid-venous, and late venous phases, and dynamic 4D-CTA [18], which continuously or intermittently acquires images throughout the passage of contrast to provide detailed temporal vascular information. Despite these advances, single-phase CTA (which requires extensive post-processing to remove background bone/soft tissue and to distinguish arteries from veins) remains the predominant clinical protocol. Robust, automated methods the brain vasculature in single-phase CTA could thus have substantial clinical impact, either by directly assisting radiologists during visual inspection of scans or by serving as an auxiliary input to pipelines for tasks such as occlusion localization [13, 16] and aneurysm detection [5, 9, 15].
Most efforts toward the segmentation of brain vessels in CTA have centered on the the automatic segmentation of the Circle of Willis (CoW), an important arterial structure. The TopCoW dataset [27] includes 125 CTA scans which were annotated using a manual, VR-based 3D workflow. The dataset has fine-grained labels for the arteries in the CoW and has drawn significant scientific interest [4, 20]. Nevertheless, focusing only on the CoW ignores parts of the brain vasculature that are relevant for the treatment of various cerebrovascular pathologies. As a result, two incremental efforts have sprung from TopCoW: a) TopBrain [1], which expands the labels in a subset of the scans from TopCoW to include veins and arteries; and b) VesselVerse, which augments the labels for all of TopCoW with manual and automatic annotations of more arteries and some veins (albeit collapsed into a single vessel class). On the other hand, other researchers have sought to address the limited availability of annotated scans by training deep learning segmentation models on large amounts of data from various domains. One such example is VesselFM [26], which relies on annotations from many domains (CTA, MRA, XR, microscopy, etc.) combined with a huge number of synthetic images. However, the fraction of CTA images they train on is small (under 1%) and no evaluations were done for head CTA vessel segmentation.
@@
Considering the above, there is a clear need for new annotation methodologies that minimize the time required to produce fine-grained labels of vessels in brain CTA. However, accurate annotation of CTA images requires specialized anatomical knowledge due to the presence of bone, soft tissue, and non-vascular contrast-enhanced structures. To address this challenge, we use dynamic 4D-CTA, which enables near-complete removal of bone and soft tissue through baseline subtraction. The high temporal resolution of dynamic CTA also minimizes interphase motion artifacts, resulting in cleaner subtracted images and more reliable vascular depiction. These enhancements overcome the longstanding difficulties in segmenting and distinguishing intracranial arteries and veins in angiographic CT—a process typically regarded as complex and labor-intensive. Using the proposed strategy (summarized in Fig. 1), we annotate a dataset comprising scans from 39 patients with artery and vein labels, and we use it to train an nnUNet [10] model, achieving remarkable artery and vein segmentation performance across contrast phases. In summary, our contributions are as follows:
@@
  • We present DynaVessel, an intracranial vessel segmentation dataset comprising 277 head CTA scans from 41 patients with artery and vein annotations
@@ 2.1 Data
This retrospective study received IRB approval with a waiver of informed consent. The 4D-CTA scans obtained from our clinical research database were acquired using a 320-detector-row CT scanner (Canon-Toshiba Aquilion One, Toshiba Medical Systems, Tokyo, Japan). Imaging was conducted in axial mode using the following parameters: 80 kV, 150 mA, 0.75 s rotation time, 320 slices with 0.5 mm thickness, and a 512 × 512 matrix. For each patient, image acquisition began roughly 7 seconds after injecting 75–100 mL of iodinated contrast agent (Omnipaque 350, General Electric Healthcare, Chicago, USA) at a rate of 4–5 mL/s, continuing intermittently through the baseline, arterial, and venous phases. As shown in Fig. 2, images were captured at 19 time points over a 60-second interval. The scanner uses the baseline image—acquired before contrast reaches the brain—to subtract bone and soft tissue from subsequent images.
Figure 3: Pre-processing steps for a dynamic CTA from a single patient. For both phases, a head ROI is used on the CTA and the subtracted CTA to suppress signals from outside the skull. Next, the subtracted images are processed using Algorithm 1 to create the vessel-separated images.
Of 106 patients scanned in 2023, 38 had scans with at least two phases of acceptable image quality. After reviewing the arterial and venous phase images and excluding those with artifacts, data from 27 patients was annotated and used for model training. For evaluation, we retrieved data from 1910 patients scanned between 2016 to 2022, of whom 140 had dynamic CTA with at least two phases of acceptable resolution. From these, 68 patients had both arterial and venous phase images, each including the original post-contrast CT image and the bone/soft tissue-subtracted image, resulting in 4 or more images per patient, covering the different CTA phases. A subset of 14 patients were randomly selected to evaluate the performance of deep learning (DL) models trained using our dataset.
@@ 2.1.2 Ground truth (GT) segmentation
For the evaluation dataset, 20 arteries were annotated using iCafe: ICA (Left), ICA (Right), VA (Left), VA (Right), BA, PComm (Left), PComm (Right), P1 (Left), P1 (Right), P2+ (Left), P2+ (Right), M1 (Left), M1 (Right), M2+ (Left), M2+ (Right), A1 (Left), A1 (Right), A2+ (Left), A2+(Right), and Acomm. A subspecialty-trained neuroradiologist validated the annotations by overlaying the segmentations with the subtracted images $S_{a}$ in 3D Slicer [14] and corrected inaccuracies related to anatomical variations in the arteries. iCafe lacks a dedicated venous annotation pipeline, so the GT vessel segmentations of the veins for the evaluation dataset were manually refined by using 3D Slicer to remove distal branches and isolate the cavernous sinuses from the overall venous segmentation. Fig. 4 depicts segmentations from our train (d) and test (e) data, compared with three other datasets (a, b, c).
@@ 2.2 Deep learning model training
Taking advantage of dynamic CTA images, our training and evaluation data includes multiple CTA scans from the same patient at different time points, each paired with a corresponding GT segmentation. All images were rigidly registered to the arterial phase CTA using Advanced Normalization Tools [23]. We split the training and test datasets at the patient level, to prevent a single patient from appearing in both validation and training sets. To train the nnUNet model, we use the nnUNet v2 library [10, 11]. Aside from specifying the model type to be a Residual Encoder UNet (large preset), all models were trained for 1000 epochs with a linearly decaying learning rate starting with 0.01, using the nnUNet planner’s automatically selected hyper-parameters. All model training and evaluation was done on a workstation with an NVIDIA RTX 4090 GPU (24 GB VRAM), 8 CPU cores, and 64 GB of RAM.
Dataset Num. train Num. test Labels Coverage
Contr. Pr. 1O 2O Di. Br. V.
VesselFM [26] 510,902 Vessel-only
TopCoW [27] 112 13 13 arteries
ISLES [19, 27] 26 13 arteries
VesselVerse [8] 112 13 Vessel-only
TopBrain [1, 27] 25 6 veins / 34 arteries
DynaVessel train 112 Vein / artery
DynaVessel1 14 Vein / 22 arteries
DynaVessel+ 165 Vein / 22 arteries
@@ 2.3 Evaluation
    Average directed Hausdorff surface distance: this variation of the Hausdorff distance was chosen based on previous work  [2, 3]. $adHD$ is similar to the balanced Hausdorff distance, but we use only the first term in the equation. This is because, in our application, the boundary of the predicted volume can be significantly larger than the GT. Thus, $adHD$ is defined as:
@@
3 Results
To assess the usefulness of DynaVessel, we compare it to two publicly available CTA vessel segmentation datasets of similar size: TopCoW [27], which contains 125 scans with semantic annotations of the arteries in the Circle of Willis; and VesselVerse [8], which expands the annotations in TopCoW to include veins and more arteries, although it uses a single vessel label. We thus train three nnUNet models as specified in Section 2.2 (from now on, we refer to the model trained on DynaVessel as $M_{DV}$). Likewise, we compare against the pre-trained VesselFM [26] DynUNet, a foundation vessel segmentation model optimized on a collection of 928 images (including TopCoW) from vessel segmentation datasets encompassing various domains, 10,000 images sampled from a generative flow-matching model, and 500,000 randomly generated 3D images imitating blood vessels based on graph representations of corrosion casts. Similar to VesselVerse, the labels in VesselFM are binary.
@@ 3 Results
Figure 5: mDC and adHD scores across contrast phases for our model, $M_{DV}$.
We evaluate all four models on several test datasets, including held-out subsets of DynaVessel, TopCoW [27], and VesselVerse [8]. We also use a subset of ISLES 2024 [19] annotated by the organizers of the TopCoW challenge [27], and TopBrain, a subset of TopCoW with detailed annotations of 34 arteries and 6 venous structures. Since DynaVessel includes images captured at several contrasts for each patient, we define two versions of it for evaluation: DynaVessel1, which contains a single, arterial-phase image for each patient (14 images total), and DynaVessel+, which contains several images per patient (165 total) with all the contrast phases, allowing us to measure model robustness to changes in contrast. Table 1 offers more details about the datasets and the train/eval splits we used. We measure performance using the metrics defined in Section 2.3 in three main ways, namely: a) the binary vessel segmentation setting, where classes in the GTs and the predictions are merged into 0 = background, 1 = vessel; b) the artery vs vein segmentation setting, where classes in the GTs are merged into 0 = background, 1 = artery, 2 = vein; and c) the semantic segmentation setting, where we evaluate performance on all available arterial anatomical locations.
@@
Figure 7: mDC scores for the semantic segmentation setting. For TopBrain, annotations of right/left distal arteries and branches are merged together to improve readability. We omit the early arterial phase images when calculating mDC on DynaVessel.
Figure 8: Qualitative comparisons of the segmentations produced by several models. In the top row, we display results for artery segmentation using (a) VesselFM, (b) the nnUNet trained on VesselVerse, and (c) our $M_{DV}$ trained on DynaVessel; true positive segmented areas are colored red, false positives in gray, and false negatives in green. The bottom row includes the results for vein segmentation using (d) VesselFM, (e) the nnUNet trained on VesselVerse, and (f) our $M_{DV}$ trained on DynaVessel; true positive segmented areas are colored in blue, false positives in gray, and false negatives in green. We again note that VesselFM and the VesselVerse nnUNet only output a single vessel class, so arterial predictions show up in the vein segmentation visualization as false positives, and vice-versa. Visualizations created on 3D Slicer.
@@
In Table 2, we present the results for binary vessel segmentation. Despite being trained on a similarly-sized dataset (or much smaller, when compared to VesselFM), $M_{DV}$ achieves higher performance on the three external test datasets and much higher performance on both internal test datasets. Remarkably, training on DynaVessel leads to better results even on data that is in-domain for the other models [^1] The TopCoW, TopBrain, and VesselVerse test sets are in-domain for VesselFM and the nnUNet models trained on TopCoW/VesselVerse, and out-of-domain for $M_{DV}$. . When the results are disaggregated my measuring performance on arteries and veins separately (see Table 3), we see that $M_{DV}$ achieves higher performance on arteries and much higher performance on veins on TopBrain and DynaVessel, underlining the positive impact of our annotation protocol across the board. The metrics for $M_{DV}$ on artery segmentation are worse on DynaVessel+ compared to the single-phase test set, which is expected as the former includes scans taken at the early arterial, venous, and late venous phases. Figs. 5 and 6 also reveal that the average scores are brought down by low performance on early arterial phases, which is expected since at that point there is likely not enough contrast liquid in the arteries.
@@
Fig. 7 shows a significant gap between the performance of $M_{DV}$ versus the two strongest baselines in the semantic segmentation evaluation, with great coverage of most anatomical regions annotated in TopBrain except for some distal arteries (the right and left anterior choroidal arteries). A similar trend is seen in the DynaVessel+ test set, where the model generally performs worse on second-order arteries compared with the first-order ones. Finally, in Fig. 8, we provide a qualitative visualization of the outputs produced by all three models. $M_{DV}$ achieves almost complete coverage of arterial and venous structures, with very few false negatives. The difference in performance is particularly noticeable on veins, where VesselFM and the VesselVerse nnUNet fail to detect most veins. Although a large number of false positive detections are visible in Fig. 8(b, c, e, f), the majority are correctly segmented arterial and venous structures that are not annotated in TopBrain. On the other hand, Fig. 8(a, d) shows that VesselFM generates a much larger number of spurious segmentations and unconnected elements, likely due to the limited amount of CTA it was trained on.

### v2

@@ Abstract
In this study, we develop a novel methodology for annotating the brain vasculature using dynamic 4D-CTA head scans. By using multiple time points from dynamic CTA acquisitions, we subtract bone and soft tissue to enhance the visualization of arteries and veins, reducing the effort required to obtain manual annotations of brain vessels. We then train deep learning models on our ground truth annotations by using the same segmentation for multiple phases from the dynamic 4D-CTA collection, effectively enlarging our dataset by 4 to 5 times and inducing robustness to contrast phases. In total, our dataset comprises 110 training images from 29 patients and 165 test images from 14 patients. In comparison with two similarly-sized datasets for CTA-based brain vessel segmentation, a nnUNet model trained on our dataset can achieve significantly better segmentations across all vascular regions, with an average mDC of 0.846 for arteries and 0.957 for veins in the TopBrain dataset. Furthermore, metrics such as average directed Hausdorff distance (adHD) and topology sensitivity (tSens) reflected similar trends: using our dataset resulted in low error margins (aDHD of 0.304 mm for arteries and 0.078 for veins) and high sensitivity (tSens of 0.877 for arteries and 0.974 for veins), indicating excellent accuracy in capturing vessel morphology. Our code and model weights are available online.
@@ 1 Introduction
Intracranial computed tomographic angiography (CTA) is an essential diagnostic tool that employs intravenous contrast agents to enhance vascular visualization, aiding in the detection of intracranial vascular abnormalities [22]. Since its introduction, CTA has advanced beyond single-phase acquisitions to include multiphase CTA [7], which captures arterial, mid-venous, and late venous phases, and dynamic 4D-CTA [16], which continuously or intermittently acquires images throughout the passage of contrast to provide detailed temporal vascular information. Despite these advances, single-phase CTA (which requires extensive post-processing to remove background bone/soft tissue and to distinguish arteries from veins) remains the predominant clinical protocol. Robust, automated methods to extract the brain vasculature in single-phase CTA could thus have substantial clinical impact, either by directly assisting radiologists during interpretation of scans or by serving as an auxiliary input to pipelines for tasks such as occlusion localization [11, 14] and aneurysm detection [5, 13].
Most efforts for brain vessel segmentation in CTA have centered on the automatic segmentation of the Circle of Willis (CoW), an important arterial structure. The TopCoW dataset [25] includes 125 CTA scans which were annotated using a manual, VR-based 3D workflow. The dataset has fine-grained labels for the arteries in the CoW and has drawn significant scientific interest [4, 18]. Nevertheless, focusing only on the CoW ignores parts of the brain vasculature that are relevant for the management of various cerebrovascular diseases. As a result, two incremental efforts have sprung from TopCoW: a) TopBrain [1], which expands the labels in a subset of the scans from TopCoW to include both veins and arteries; and b) VesselVerse [2], which augments the labels for all of TopCoW with manual and automatic annotations of more arteries and some veins (albeit collapsed into a single vessel class). On the other hand, other researchers have attempted to address the limited availability of annotated CTA data by training deep learning models on large, heterogeneous datasets from multiple modalities. One such example is VesselFM [24], which relies on annotations from many modalities (CTA, MRA, X-Ray, microscopy, etc.) combined with a huge number of synthetic images. However, the fraction of training CTA images is small (under 1%) and the model has not been evaluated for head CTA vessel segmentation.
@@
Considering the above, there is a clear need for new annotation methodologies that minimize the time required to produce fine-grained labels for vessels in brain CTA. However, accurate annotation of CTA images requires specialized anatomical knowledge due to the presence of bone, soft tissue, and non-vascular contrast-enhanced structures. To address this challenge, we use dynamic 4D-CTA, which enables near-complete removal of bone and soft tissue through baseline subtraction. The high temporal resolution of dynamic CTA also minimizes interphase motion artifacts, resulting in cleaner subtracted images and more reliable vascular depiction. These enhancements overcome the longstanding difficulties in segmenting and distinguishing intracranial arteries and veins in angiographic CT—a process typically regarded as complex and labor-intensive. Using the proposed strategy (summarized in Fig. 1), we annotate a dataset comprising scans from 39 patients with artery and vein labels, and we use it to train an nnUNet [8] model, achieving remarkable artery and vein segmentation performance across contrast phases. In summary, our contributions are as follows:
@@
  • We present DynaVessel, an intracranial vessel segmentation dataset comprising 275 head CTA scans from 43 patients with artery and vein annotations
@@ 2.1 Data
This retrospective study received IRB approval with a waiver of informed consent. The 4D-CTA scans obtained from our clinical research database were acquired using a 320-detector-row CT scanner (Canon-Toshiba Aquilion One, Toshiba Medical Systems, Tokyo, Japan). Imaging was conducted in axial mode using the following parameters: 80 kV, 150 mA, 0.75 s rotation time, 320 slices with 0.5 mm thickness, and a 512 × 512 matrix. For each patient, image acquisition began roughly 7 seconds after injecting 75–100 mL of iodinated contrast agent (Omnipaque 350, General Electric Healthcare, Chicago, USA) at a rate of 4–5 mL/s, continuing intermittently through the baseline, arterial, and venous phases. As shown in Fig. 2, images were captured at 19 time points over a 60-second interval. The peak arterial and venous phase images, which correspond to $t=27.6$s and $t=45.3$s respectively, are of primary interest. The scanner uses the baseline image—acquired before contrast reaches the brain—to subtract bone and soft tissue from subsequent images.
Figure 3: Pre-processing steps for a dynamic CTA study. For both phases, a head ROI is used on the CTA and the subtracted CTA to suppress the background. Next, the subtracted images are processed using Algorithm 1 to create the vessel-separated images.
Of the 106 patients scanned in 2023, 38 had scans containing at least two phases of acceptable image quality. Although these were acquired using dynamic CTA protocols, complete phase datasets were not consistently archived in the PACS; in many cases, only thicker slice reconstructions, subtracted volumes, or MIP images were retained. After review, 29 patients were identified as having artifact-free arterial and venous phase images. These scans were annotated and used for model training. For evaluation, we retrieved data from 1910 patients scanned between 2016 to 2022, of whom 140 had dynamic CTA with at least two phases of acceptable resolution. Among these, 68 patients had both peak arterial and venous phase images (each comprising the original post-contrast CT image and the corresponding bone/soft tissue-subtracted image), with many also including more CTA phases. A subset of 14 patients were randomly selected to evaluate the performance of deep learning (DL) models trained using our dataset.
Dataset Num. train Num. test Labels Coverage
Contr. Pr. 1O 2O Di. Br. V.
VesselFM [24] 510,902 Vessel-only
TopCoW [25] 112 13 13 arteries
ISLES [17, 25] 26 13 arteries
VesselVerse [2] 112 13 Vessel-only
TopBrain [1, 25] 25 6 veins / 34 arteries
DynaVessel train 110 Vein / artery
DynaVessel1 14 Vein / 22 arteries
DynaVessel+ 165 Vein / 22 arteries
@@ 2.1.2 Ground truth (GT) segmentation
For the evaluation dataset, 20 arteries were annotated using iCafe: ICA (Left), ICA (Right), VA (Left), VA (Right), BA, PComm (Left), PComm (Right), P1 (Left), P1 (Right), P2+ (Left), P2+ (Right), M1 (Left), M1 (Right), M2+ (Left), M2+ (Right), A1 (Left), A1 (Right), A2+ (Left), A2+(Right), and Acomm. A subspecialty-trained neuroradiologist validated the annotations by overlaying the segmentations with the subtracted images $S_{a}$ in 3D Slicer [12] and corrected inaccuracies related to anatomical variations in the arteries. iCafe lacks a dedicated venous annotation pipeline, so the GT vessel segmentations of the veins for the evaluation dataset were manually refined by using 3D Slicer to remove distal branches from the venous segmentation. Fig. 4 depicts segmentations from our train (d) and test (e) data, compared with three other datasets (a, b, c).
@@ 2.2 Deep learning model training
Taking advantage of dynamic CTA images, our training and evaluation data includes multiple CTA phases from the same patient, each paired with a corresponding GT segmentation. All images were rigidly registered to the arterial phase CTA using Advanced Normalization Tools [21]. We split the training and test datasets at the patient level, to prevent any patient from appearing in both validation and training sets. To train the nnUNet model, we use the nnUNet v2 library [8, 9]. Aside from specifying the model type to be a Residual Encoder UNet (large preset), all models were trained for 1000 epochs with a linearly decaying learning rate starting at 0.01, using the nnUNet planner’s automatically selected hyper-parameters. All model training and evaluations were performed on a workstation with an NVIDIA RTX 4090 GPU (24 GB VRAM), 8 CPU cores, and 64 GB of RAM.
@@ 2.3 Evaluation
    Average directed Hausdorff surface distance: this variation of the Hausdorff distance was chosen based on previous work  [3]. $adHD$ is similar to the balanced Hausdorff distance, but we use only the first term in the equation. This is because, in our application, the boundary of the predicted volume can be significantly larger than the GT. Thus, $adHD$ is defined as:
@@
Figure 6: mDC scores for semantic segmentation. For TopBrain, we merge annotations of right/left distal arteries and branches to improve readability. We omit the pre-contrast/early arterial phase images when calculating mDC on DynaVessel.
@@
3 Results
To assess the usefulness of DynaVessel, we compare it to two publicly available CTA vessel segmentation datasets of similar size: TopCoW [25], which contains 125 scans with semantic annotations of the arteries in the Circle of Willis; and VesselVerse [2], which expands the annotations in TopCoW to include veins and more arteries, although it uses a single vessel label. We thus train three nnUNet models as specified in Section 2.2 (from now on, we refer to the model trained on DynaVessel as $M_{DV}$). Likewise, we compare against the pre-trained VesselFM [24] DynUNet, a foundation vessel segmentation model optimized on a collection of 928 images (including TopCoW) from vessel segmentation datasets encompassing various modalities, 10,000 images sampled from a generative flow-matching model, and 500,000 randomly generated 3D images imitating blood vessels based on graph representations of corrosion casts. Similar to VesselVerse, the labels in VesselFM are binary.
Figure 7: Qualitative comparisons of the segmentations from several models on a TopBrain sample. In the top row, we display results for artery segmentation using (a) VesselFM, (b) the VesselVerse nnUNet, and (c) our $M_{DV}$; true positive (TP) areas are red, false positives (FPs) are gray, and false negatives (FNs) are green. The bottom row includes the results for vein segmentation using (d) VesselFM, (e) the VesselVerse nnUNet, and (f) our $M_{DV}$; TP areas are blue, FPs are gray, and FNs are green. VesselFM and the VesselVerse nnUNet only output a single vessel class, so arterial predictions show up in the vein segmentation visualization as FPs, and vice-versa. We note that most FP areas segmented by $M_{DV}$ (c, f) are actually TP arterial and venous structures that are not annotated in TopBrain. Visualizations created on 3D Slicer.
We evaluate all four models on several test datasets, including held-out subsets of DynaVessel, TopCoW [25], and VesselVerse [2]. In addition, we use a subset of ISLES 2024 [17] annotated by the organizers of the TopCoW challenge [25], and TopBrain, a subset of TopCoW with detailed annotations of 34 arteries and 6 venous structures. Since DynaVessel includes images acquired at multiple contrasts for each patient, we define two evaluation variants: DynaVessel1, which contains a single, arterial-phase image for each patient (14 images total), and DynaVessel+, which contains all available contrast phases per patient (165 images in total), allowing us to measure model robustness to contrast variability. Table 1 offers more details about the datasets and the train/eval splits we used. Model performance is assessed using the metrics defined in Section 2.3 in three main ways, namely: a) the binary vessel segmentation setting, where classes in the GTs and the predictions are merged into 0 = background, 1 = vessel; b) the artery vs vein segmentation setting, where classes in the GTs are merged into 0 = background, 1 = artery, 2 = vein; and c) the semantic segmentation setting, where we evaluate performance on all available arterial anatomical locations.
@@ 3 Results
In Table 2, we present the results for binary vessel segmentation. Despite being trained on a similarly-sized dataset (or much smaller, when compared to VesselFM), $M_{DV}$ achieves higher performance on the three datasets. Remarkably, training on DynaVessel leads to better results even on data that is in-domain for the other models [^1] The TopCoW, TopBrain, and VesselVerse test sets are in-domain for VesselFM and the nnUNet models trained on TopCoW/VesselVerse, and out-of-domain for $M_{DV}$. . When the results are disaggregated to arteries and veins separately (see Table 3), we see that $M_{DV}$ achieves higher performance on arteries and much higher performance on veins on TopBrain and DynaVessel, underlining the positive impact of our annotation protocol across the board. The metrics for $M_{DV}$ on artery segmentation are worse on DynaVessel+ compared to the single-phase test set, which is expected as the former includes scans taken before contrast and at the early arterial, venous, and late venous phases. Fig. 5 also reveals that the average scores are brought down by low performance on pre-contrast and early arterial phases, which is expected since at that point there is likely not enough contrast in the arteries.
@@
Fig. 6 shows a significant gap between the performance of $M_{DV}$ versus the two strongest baselines in the semantic segmentation evaluation, with great coverage of most anatomical regions annotated in TopBrain except for some distal arteries (the right and left anterior choroidal arteries). A similar trend is seen in the DynaVessel+ test set, where the model generally performs worse on second-order arteries compared with the first-order ones. Meanwhile, in Fig. 7, we provide a qualitative visualization of the outputs produced by all three models. $M_{DV}$ achieves almost complete coverage of arterial and venous structures, with very few false negatives. The difference in performance is particularly noticeable on veins, where VesselFM and the VesselVerse nnUNet fail to detect most veins. In VesselVerse’s case, this is due to the fact venous sinuses were not annotated to maintain consistency with the MRA annotations. Although a large number of false positive detections are visible in Fig. 7(b, c, e, f), the majority are correctly segmented arterial and venous structures that are not annotated in TopBrain. On the other hand, Fig. 7(a, d) shows that VesselFM generates a much larger number of spurious segmentations and unconnected elements, likely due to the limited amount of CTA it was trained on.
3.1 Ablation study
In Fig. 8, we measure the impact of removing the various contrast phases used for training $M_{DV}$ by optimizing two new models: $M_{art}$, which is trained on only 29 peak arterial phase images, and $M_{ven}$, which is trained on only 29 peak venous phase images. It is important to note that $M_{art}$ and $M_{ven}$ have access to the same number of patients and GT annotations as $M_{DV}$. We make two observations: 1) that training on a single, peak phase per patient significantly reduces the models’ capabilities to generalize across unseen phases; and 2) that the model trained on all available phases often surpasses the models trained on a single phase when evaluated on such phases. This once again confirms the usefulness of our methodology, especially in scenarios where the availability of peak arterial/venous phase images is not guaranteed.
Figure 8: Artery and vein mDC scores across contrast phases for our model ($M_{DV}$, violet, rightmost box in each subplot), compared with models trained either on the peak arterial phase images ($M_{art}$, red, leftmost on each box) or the peak venous phase images ($M_{ven}$, blue, middle box in each subplot). The number above the top whisker for each boxplot codifies the difference between the average Hounsfield unit (HU) intensity in arteries/veins during training vs during the phase on which the model is evaluated. The performance for single-phase training correlates negatively with the difference between mean training HU and mean test HU. The behavior in the late venous phase is explained by the circulation of part of the contrast agent back into arteries.